OpenStudy (anonymous):

Find the value of c if q = 9600 joule, m =120 g and Ti = 25.0 oC and Tf = 36.2 oC

OpenStudy (souvik):

what is c?

OpenStudy (aaronq):

q=m*C*dT=m*C*(Tf-Ti)